OpenStudy (anonymous):

Find all radian solutions: sinx + 1 = cosx

OpenStudy (anonymous):

sin^2 x + cos^2 x = 1 sin x = sqrt (1-cos^2 x)

OpenStudy (anonymous):

sqrt (1-cos^2 x) = (cosx - 1) squaring both sides 1- cos^2 x = cos^2 x - 2cosx + 1 2 cos^2 x -2cosx = 0 hope nw u can solve